OpenStudy (anonymous):

Find all the eighth roots of (19 + 7 i)

OpenStudy (anonymous):

19 + 7 i = √410 [ cos (0.35299 + 2nπ) + i sin (0.34299) ] ( 19 + 7 i)^(1/8) = (410)^(1/16) [ cos ((0.35299 + 2nπ) / 8 ) + i sin (( 0.35299 + 2 n π) / 8 ) ] for n= 0,1,2,3,4,5,6,7....you compute